OpenStudy (anonymous):

A country's population in 1992 was 50 million. In 2002 it was 53 million. Estimate the population in 2010 using the exponential growth formula

OpenStudy (tennistar):

It would be about 55 million, because 3 million is added every ten years. And you have 8 years

OpenStudy (anonymous):

Exponential growth formula\[P(t) = Ae^{kt}\] \[P(0) = A = 50000000\]\[P(2002 - 1992) = P(10) = 50000000e^{10k} = 53000000\]\[k = 0.058269\]\[P(2010 - 1992) = P(18) = 50000000e^{18(0.0058269)} = 55 529 091\]

OpenStudy (anonymous):

oops k = 0.0058269